On January 1, 2021, Mania Enterprises issued 15% bonds dated January 1, 2021, with a face amount of $28 million. The bonds mature in 2033 (10 years). For bonds of similar risk and maturity, the market yield is 12%. Interest is paid semiannually on June 30 and December 31. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) Required: Determine the price of the bonds at January 1, 2021. (Enter your answer in whole dollars.)

Price of Bonds:
Answer #1
(note there will be slight difference in issue price calculation ,kindly use the factors as given
inn your question to get exact answer)
Bond characterstics Amount
1-a) Principal 28,000,000
interest        2,100,000
Market interest rate 6%
periods to maturity 20
issue price 32,817,232 answer
Calculation of bond issue price
Where
i= 6.00%
t= 20
principal * PV of $1 at 6% for 20 yrs =
28,000,000 * 0.3118        = 8730400
interest * PV of ordinary annuity at 6%=
2100000 * 11.46992 = 24086832
bond issue price 32817232